Hello, My Name is Donna and I must say, I am so pleased you have the ratings in the choices viewed. Believe me when I say I truly need assistance with my Math 126 Class. I am eager to learn however as I love the expansion of knowledge at any time!
My question relates to the working of and creating the equasion I will copy and paste. I am a week behind in class due to MS and Arthritis. Been working with fervor to get caught up.I thank you for any and all assistance you may be able to provide me.
Best Regards,
Donna Vie

~~~~~~~~~~~~~~~~~~~~~~~~


2. In years, how old was the older person when you were born? 8
3. Write an equation that models how old in years each of you will be, when your ages add up to 150 years old. For example, if x = your age and the eldest person was a year older than you, you would write their age as x + 1. Then the equation would be: x + (x+1) = 150.
X + (x + 8) = 150
I can not seem to grasp this fully. My numbers do not come out to 150 and I cannot see what I am doing wrong? Help anyone??
4. Explain the reasoning which helped you develop your equation.
5. Solve for your future ages. Are your answers reasonable, do they add up to 150?
6. In years, how old were you when the youngest person was born?
7. At some point during the lives of you and the youngest person, your age will be three times his/her age at that moment. Write an equation which models how old in years each of you will be when you are three times as old as the younger person.
8. Explain the reasoning which helped you develop your equation.
9. Solve the equation for your ages when you are three times as old as the youngest person. Are your answers reasonable?

Answer by dragonwalker(72)

Well, if the older person was 8 when you were born and you want to know how old you are when both ages add up to 150:
You are correct in your equation:
x + (x + 8) = 150
you need to rearrange to find x:
so you put the x's on one side and numbers on the other. Remember if you move something to the other side of the equation you change the sign (unless you are just switching everything)
So:
x + x + 8 = 150
x + x = 150 - 8
2x = 142
to find x you need to divide both sides by 2 to change the 2x to x:
2x/2 = 142/2
x = 71
so you are 71 and the older person is 71+8 = 79


For an equation stating that the older person is 3 times the younger person. If you are still using the same i.e. x + 8 for the older person:
Then 3x = x + 8
both sides of the equation represent the age of the older person. At the same time they are three times your age but they are still 8 years older at the same time so both equations (3x and x+8) are equal to each other.
so if 3x = x + 8
rearrange as before:
3x - x = 8
2x = 8
2x/2 = 8/2
x = 4
and they would be x + 8 i.e. 4 + 8 = 12